A gas occupies a volume of 50 mL at a temperature of 0°C and a pressure of 2,000 N / m3. What volume will be occupied by the same amount of gas also at 0°C if the pressure is increased to 10,000 N / m3.


1
Expert's answer
2022-02-06T14:32:41-0500

Given:

"V_1=\rm 50\: mL"

"p_1=\rm 2000\: N/m^2"

"p_2=\rm 10000\: N/m^2"

"V_2-?"


The Boyle–Mariotte law says

"p_1V_1=p_2V_2"

Hence

"V_2=V_2\frac{p_1}{p_2}"

"V_2=50\:\rm mL*\frac{2000\: N/m^2}{10000\: N/m^2}=10\: mL"